### v2

# One-Shot Imitation under Mismatched Execution

Abstract
Human demonstrations as prompts are a powerful way to program robots to do long-horizon manipulation tasks. However, translating these demonstrations into robot-executable actions presents significant challenges due to execution mismatches in movement styles and physical capabilities. Existing methods either depend on robot-demonstrator paired data, which is infeasible to scale, or rely too heavily on frame-level visual similarities that often break down in practice. To address these challenges, we propose RHyME, a novel framework that automatically aligns robot and demonstrator task executions using optimal transport costs. Given long-horizon robot demonstrations, RHyME synthesizes semantically equivalent demonstrator videos by retrieving and composing short-horizon demonstrator clips. This approach facilitates effective policy training without the need for paired data. We demonstrate that RHyME outperforms a range of baselines across cross-embodiment datasets, showing a 52% increase in task recall over prior cross-embodiment learning methods. We release our code and datasets at website.
Figure 1: Overview of RHyME. We introduce RHyME, a hierarchical framework that trains a robot policy to mimic a long-horizon video from a demonstrator that exhibits mismatched task execution. Inference Time (Left): Our robot policy translates a demonstrator video into actions to complete the same long-horizon tasks specified by the input. Train Time (Right): Given unpaired robot-demonstrator datasets, RHyME ”imagines” a paired dataset by employing sequence-level similarity metrics which can be used for training the policy.
I Introduction
Human demonstrations offer an effective approach for programming robots to execute long-horizon manipulation tasks [1, 2, 3, 4]. Unlike language instructions, demonstrations are grounded in the task environment, providing rich with cues for what steps to follow, which objects to interact with, and how to interact with them [5, 6].
We view this as a translation problem where a human video must be translated into a series of robot actions [7, 8, 9, 10]. However, training such policies typically requires paired human-robot demonstrations, which is impractical to collect at scale for long-horizon tasks. Although large-scale human demonstration videos (e.g., YouTube) and robot datasets exist [11, 12], they are unpaired, making them unsuitable for directly learning this translation.
Prior works leverage unpaired human and robot demonstrations to learn visual representations that map both human and robot images into a shared embedding space [4, 6, 13, 14, 15]. A policy is then trained to generate actions conditioned on robot video embeddings, and directly transferred at test time to work with embeddings from a human prompt video. However, a key assumption these works rely on is that the demonstrator and robot perform tasks with matched execution, i.e., the human executes tasks in a visually similar way to that of the robot (e.g. slowly moving one arm with a simple grasp). In reality, humans often act more swiftly, use both hands for manipulation, or even execute multiple tasks simultaneously, creating a mismatch in execution styles. This mismatch leads to misalignment between the demonstrator and robot embeddings, hindering direct policy transfer.
We tackle this problem of imitation under mismatched execution. Our key insight is that while a robot and demonstrator may perform the same task in visually and physically different ways, we can establish a high-level equivalence by reasoning over the entire sequence of image embeddings they generate. We show that while individual image embeddings may appear different between robot and demonstrator, we can construct sequence-level similarity functions where the two are closer. Notably, we can do this without fine-tuning representations on paired data.
We propose RHyME (Retrieval for Hybrid Imitation under Mismatched Execution), a framework which trains a robot policy to follow a long-horizon demonstration from a mismatched expert without access to paired robot-demonstration videos (Fig. 1). First, RHyME defines a sequence-level similarity metric between robot and demonstrator embeddings, using optimal transport to measure alignment. Given a robot trajectory and a database of demonstrator data, RHyME imagines a long-horizon video by retrieving and composing short-horizon snippets of demonstrations similar to the robot video. This retrieval process is guided by an optimal transport similarity metric between robot and demonstrator sequences. The framework trains a policy using a hybrid approach, incorporating both real robot demonstrations and imagined demonstrator sequences. Our contributions are:
  1. 1.
    We propose RHyME, a novel retrieval-based algorithm for one-shot initiation from videos of a demonstrator with mismatched execution. Without access to paired datasets, RHyME utilizes high-level equivalence between robot and demonstrator videos using optimal transport costs.
  2. 2.
    We systematically study the performance of different video retrieval by proposing three novel cross-embodiment datasets in the Franka Kitchen simulator. The demonstration datasets, totaling 10 hours of videos, are designed to vary in visual and execution mismatches relative to the robot.
  3. 3.
    We show that RHyMe outperforms a range of baselines on all degrees of mismatches, yielding over 50% increase in success rates in the most challenging scenarios.
II Related Work
Demonstrations have been utilized to guide robot manipulation policies in several different ways. We place our work relative to each cluster of related research.
Tracking Reference Motion. The imitation challenge is reduced to motion tracking when the robot receives the demonstrator’s motion as input. Robots with human-like joint configurations can directly mimic human trajectories, a method applied to humanoid robots with similar morphologies [16, 17, 18, 19, 20, 21] and robotic hands [22, 23, 24, 25, 26, 27, 28, 29]. When the execution capabilities are mismatched, either human models are simplified [16, 18, 20] or robot trajectories [17, 19, 30, 31] are optimized to approximately match the reference trajectory. [21] shows a method to extract human poses from videos and use that as a reference for training reinforcement learning agents. Similarly, human demonstrations have been used to guide reinforcement learning for robotic hands [22, 23, 24, 25]. More recently, large-scale video datasets of humans on the internet have been used to extract hand positions and adapted for robot manipulation [28, 27, 29, 26, 32], or rely on other common abstractions such as optical flow as a common trajectory representation across embodiments [33], still requiring execution to be matched. Distinct from these works, we focus on learning robot manipulation tasks directly from demonstrators’ RGB videos without explicit motion input.
Learning Reward Functions from Demonstrator Videos. These methods tackle the problem of matching the demonstrator when their motion cannot be simply mimicked. Still, their videos contain useful task information and can be used to learn reward functions for reinforcement learning on the robot. A general method in this line of research [34, 35, 36, 2] is to extract reward functions that encourage the robot to manipulate objects in the same way as the video. For example, [34] self-supervises the robot by making it learn to match the demonstrations of a demonstrator perturbing a rope. GraphIRL [2] extracts sequences of object pose movement from the demonstrator and enforces temporal cyclic consistency with the robot’s execution. Other approaches frame the problem as task matching [37, 38], where the robot is rewarded when it is deemed to perform the same skill as the demonstrator. While these papers tackle the problem of mismatched execution directly, they require reinforcement learning to train robot policies, which can be challenging for complex tasks.
Learning Aligned Robot-Demonstrator Representations. Another strategy for addressing this challenge is to train representations for both the robot and the demonstrator that are indistinguishable when performing the same task. This approach often frames the task as a video translation challenge, where the demonstrator’s video is converted into a robot’s perspective to simplify task mimicry [10, 9, 39]. Additionally, methods like WHIRL [3] align videos by effectively masking out both the robot and the demonstrator, creating a neutral visual field. Embeddings can be aligned using datasets that either directly pair human and demonstrator actions or utilize human preference datasets to rank image frames, as shown in X-IRL [1] and RAPL [40]. Unlike these methods, our approach does not depend on labeled correspondences between robots and demonstrators.
One-shot Visual Imitation from Demonstration Videos. We tackle this problem setting in our work where the robot imitates actions from demonstrator videos in a one-shot setting, i.e., the robot uses a prompt video as a guide, aiming to replicate the demonstrated actions after viewing them once [7, 8, 4, 6, 41, 42, 43, 32]. If a paired dataset of robot and demonstrator videos executing the same task exists, the robot can learn to translate a prompt video into actions directly [7, 8]. The closest to our work is the setting without paired data of demonstrator and robot skills. Prior works [6, 4] train policies conditioning on robot videos and zero-shot transfer to a prompt demonstration at test time using aligned visual embeddings. For example, XSkill [4] uses a self-supervised clustering algorithm based on visual similarity to align representations of demonstrations and robot videos. However, such an approach can falter when there are significant mismatches in execution. We address this issue by posing the visual imitation problem as one of train-time retrieval. During training, we match robot videos to the closest demo videos from an unpaired play dataset to imagine synthetic demo videos. Training robot policies conditioned on these synthetic videos enable the robot to translate demonstration videos into robot actions at test time.
III Problem Formulation
Inference Time: Translate Demonstration Video to Robot Actions. The robot’s goal is to replicate a series of tasks demonstrated in a video using a policy $\pi(a_{R}|s_{R},\mathbf{v_{H}})$ that translates the video into robot actions $a_{R}$ at state $s_{R}$. The demonstration video is a sequence of images $\mathbf{v_{H}}=\{v_{H}^{0},v_{H}^{1},\dots,v_{H}^{T}\}$, where $T$ is the length of the video.
Train Time: Learning from Unpaired Demonstrator and Robot Data. While training a policy with paired robot and demonstrator data is feasible, collecting such data at scale is impractical. Instead, we frame the problem as learning and leveraging aligned embeddings from unpaired data, enabling the transfer of policies trained on robot embeddings to demonstration embeddings.
We assume access to two datasets —a robot dataset ($D_{\rm robot}$) of long-horizon manipulation tasks and a play dataset ($D_{\rm play}$) of short-horizon demonstrator video clips showing interactions with objects and the environment. The robot dataset, $D_{\rm robot}=\{(\mathbf{\xi_{R}},\mathbf{v_{R}})\}$, comprises pairs of state-action trajectories and robot videos. Each robot trajectory, $\mathbf{\xi_{R}}=\{(s_{R}^{0},a_{R}^{0}),(s_{R}^{1},a_{R}^{1}),\dots,(s_{R}^{T},a_{R}^{T})\}$, represents the sequence of robot states and actions throughout an episode. Correspondingly, the video $\mathbf{v_{R}}=\{v_{R}^{0},v_{R}^{1},\dots,v_{R}^{T}\}$ is a sequence of images of the robot executing the task. The play dataset, $D_{\rm play}=\{\mathbf{v_{H}}\}$, consists of demonstrator videos that do not have direct correspondence with the robot dataset. At test time, the demonstrator’s video contains a set of tasks whose composition is unseen by the robot during training. However, we make an assumption that the constituent tasks are individually covered both in $D_{\rm play}$ and $D_{\rm robot}$. This assumption is consistent with prior work [4].
Our goal is to train two modules: a vision encoder and a robot policy. The vision encoder maps both robot and demonstrator videos into a shared embedding space to enable translation. We employ a video encoder $\phi(\mathbf{v})$ to extract a sequence of embeddings $\mathbf{z}=\{z_{0},z_{1},....,z_{T}\}$ for each frame from all videos [^1] We encode a 1-timestep sliding window of 8 neighboring images to generate each image embedding.. Then, given a demonstrator video $\mathbf{v_{H}}$, we generate a sequence of latent embeddings $\mathbf{z_{H}}$. We aim to train a policy that conditions on the sequence of embeddings to predict robot actions $\pi(a_{R}|s_{R},\mathbf{z_{H}})$ without access to paired robot and demonstrator data. We discuss how to train both the encoder and the policy in Section IV.
IV Approach
We present RHyME, a one-shot imitation learning algorithm that translates demonstrations to robot actions, without paired data. Before policy training, we first train a video encoder using a dataset of unpaired robot and demonstration videos (Section  IV-A). Then, this trained video-encoder is frozen and utilized for retrieval during policy training (Section IV-B). At train time, given just a robot trajectory, RHyME imagines a corresponding demonstration by retrieving and composing corresponding demonstrator snippets. It then trains a policy to predict robot actions, conditioned on the imagined demonstration. We discuss details of the retrieval, training process, and video embeddings below.
Input: Robot embeddings $\mathbf{z_{R}}$, demonstrator play dataset $D_{\rm play}$, video encoder $\phi(\mathbf{z}|\mathbf{v})$, segment length $K$, sequence-level video distance function $d$
Output: Imagined Demo $\mathbf{\hat{z}_{H}}$
Initialize empty demo $\mathbf{\hat{z}_{H}}\leftarrow\{\}$
// Divide sequence into short snippets
$Z_{R}=\{{z_{R}^{1:K}},{z_{R}^{K+1:2K}},\dots,{z_{R}^{T-K+1:T}}\}$
1for robot segment ${z_{R}^{i:i+K}}\text{ in }Z_{R}$ do
// Find closest snippet from play dataset
${{\hat{z}_{H}}}\leftarrow\operatorname*{arg\,min}\limits_{{\hat{z}_{H}}=\phi(v_{H}),v_{H}\in D_{\rm play}}{d({\hat{z}_{H}},{z_{R}^{i:i+K}}})$
// Append snippet to imagined demo
2      $\mathbf{\hat{z}_{H}}\text{.extend}({\hat{z}_{H}})$
Return Imagined demo embeddings $\mathbf{\hat{z}_{H}}$
Algorithm 1 Imagine-Demo: Retrieve and compose demonstration snippets given a robot trajectory
Input: Robot dataset ${D}_{\rm robot}$, demonstrator play dataset $D_{\rm play}$, video encoder $\phi(\mathbf{z}|\mathbf{v})$
Output: Trained robot policy $\pi_{\theta}(a|s,\mathbf{z})$
1Initialize robot policy $\pi_{\theta}$ and paired dataset $D_{\rm paired}\leftarrow\{\}$ for $(\mathbf{v_{R}},\mathbf{\xi_{R}})\text{ in }D_{\rm robot}$ do
Generate robot embeddings $\mathbf{z_{R}}=\phi(\mathbf{v_{R}})$
// Retrieve and pair demonstration
$\mathbf{\hat{z}_{H}}\leftarrow\texttt{Imagine-Demo}(\mathbf{z_{R}},{D}_{\rm play})$
// Add paired data into dataset
2      $D_{\rm paired}\leftarrow D_{\rm paired}\cup(\mathbf{\hat{z}_{H}},\mathbf{{z}_{R}},\mathbf{\xi_{R}})$
3while not converged do
// Get demo, robot embeddings with actions
$(\;\mathbf{\hat{z}_{H}},\mathbf{z_{R}},\mathbf{\xi_{R}}\;)\sim D_{\rm paired}$
// Hybrid training
4      for $(s_{t},a_{t})$ in $\mathbf{\xi_{R}}$ do
// Condition on imagined demo
$\pi_{\theta}\leftarrow\texttt{Update-Policy}(a_{t},\pi_{\theta}(s_{t},\mathbf{\hat{z}_{H}}))$
// Condition on robot video
5            $\pi_{\theta}\leftarrow\texttt{Update-Policy}(a_{t},\pi_{\theta}(s_{t},\mathbf{z_{R}}))$
6
Return Trained Robot Policy $\pi$
Algorithm 2 RHyME: Train robot policy on both robot trajectories and imagined demonstrations
IV-A Training the Vision Encoder
We align the robot and demonstrator video embeddings in three ways: visually, temporally, and at the task level, all without requiring trajectory-level correspondences. We employ unsupervised losses $\mathcal{L}_{\rm vis}(\phi)$ and $\mathcal{L}_{\rm temp}(\phi)$ for visual and temporal alignment, following prior works [44, 45], and introduce an optional task alignment loss $\mathcal{L}_{\rm task}(\phi)$.
Figure 2: Cross-Embodiment Datasets (Left): We introduce three cross-embodiment demonstrators in the Franka Kitchen simulator with varying visual and execution mismatches. Sphere-Easy replaces the robot model with a sphere at the end effector location, inducing a visual gap. Sphere-Medium additionally deviates in manipulation style by moving objects along different trajectories. Sphere-Hard resembles humans, as the demonstrator is bimanual and parallelizes tasks. Results (Right): We report performance across all datasets on a range of baselines, showing that as the demonstrator’s execution visually and physically deviate further from those of the robot, policies trained with our framework RHyME consistently outperforms XSkill.
Visual Alignment. To align robot and demonstrator embeddings ($\mathbf{z_{R}}$, $\mathbf{z_{H}}$), we use SwAV [44], a self-supervised method that clusters images based on shared visual features. SwAV learns a set of $K$ prototype vectors, to which each image is assigned. The SwAV loss $\mathcal{L}_{\rm vis}(\phi)$ updates both the encoder and prototypes, aligning robot and demonstrator videos by clustering similar visual features.
Temporal Alignment. To align temporally adjacent frames in robot and demonstrator videos, we use Time Contrastive Loss [45]. This loss encourages embeddings of frames close in time to be similar. For each frame $z^{t}$, we define a positive set $\mathbf{z^{+}}$ of frames within a temporal window $w$, and a negative set $\mathbf{z^{-}}$ for frames outside this window. Using the contrastive loss $\mathcal{L}_{\rm temp}(\phi)$, we pull embeddings from the positive set closer and push negative set embeddings further apart, capturing temporal continuity across videos.
Task Alignment. Task-level alignment $\mathcal{L}_{\rm task}(\phi)$ is used when a small set of paired robot and demonstrator snippets is available. Unlike frame-level methods, this aligns video embeddings of the robot $\mathbf{z_{R}}$ and demonstrator $\mathbf{z_{H}}$. We compute the optimal transport distance $d(\mathbf{z_{R}},\mathbf{z_{H}})$ to measure the similarity between two sequences of video embeddings. We then apply a contrastive learning objective (INFO-NCE [46]) to pull matched embeddings closer and push different-task embeddings apart. The final task alignment loss is: $\mathcal{L}_{\rm task}(\phi)=-\sum\limits_{i}\frac{\text{exp}(-d(\mathbf{z_{R}}^{i},\mathbf{z_{H}}^{i}))}{\text{exp}(-d(\mathbf{z_{R}}^{i},\mathbf{z_{H}}^{i}))+\sum_{j\neq i}\text{exp}(-d(\mathbf{z_{R}}^{i},\mathbf{z_{H}}^{j}))}$
Our final loss function for training the visual encoder $\phi$ is:
$\mathcal{L}(\phi)=\lambda_{\rm vis}\mathcal{L}_{\rm vis}(\phi)+\lambda_{\rm temp}\mathcal{L}_{\rm temp}(\phi)+\lambda_{\rm task}\mathcal{L}_{\rm task}(\phi)$ (1)
where $\lambda_{\rm task}=0$ by default and non-zero only with access to short-horizon paired data (Fig. 5).
IV-B Training the Robot Policy
Training Overview. Algorithm 2 provides an overview of our approach to train a robot policy $\pi_{\theta}$ using both robot trajectories and imagined demonstration videos. The training process has two stages.
Stage 1: Create a Paired Dataset. For each robot trajectory $\mathbf{\xi_{R}}$ and video $\mathbf{v_{R}}$ in $D_{\rm robot}$, we encode the robot video into embeddings $\mathbf{z_{R}}=\phi(\mathbf{v_{R}})$ using the learned video encoder $\phi$. We then retrieve imagined demonstration embeddings $\mathbf{\hat{z}_{H}}$ by aligning $\mathbf{z_{R}}$ with demonstration snippets from the play dataset $D_{\rm play}$, through the function Imagine-Demo. This produces a paired dataset $D_{\rm paired}$ containing $(\mathbf{\hat{z}_{H}},\mathbf{z_{R}},\mathbf{\xi_{R}})$.
Stage 2: Train Policy on Paired Dataset. The policy $\pi_{\theta}$ is trained on the paired dataset $D_{\rm paired}$ in a hybrid fashion. For each element in $D_{\rm paired}$, we update the policy in two modes —Mode 1: The policy is conditioned on the robot video embeddings $\mathbf{z_{R}}$ to predict actions $\pi_{\theta}(a_{t}|s_{t},\mathbf{z_{R}})$, Mode 2: The policy is conditioned on the imagined demonstration embeddings $\mathbf{\hat{z}_{H}}$ to predict actions $\pi_{\theta}(a_{t}|s_{t},\mathbf{\hat{z}_{H}})$. By alternating between these two modes, the policy learns to generalize from both robot and imagined demonstration videos, enabling it to handle execution mismatches.
Figure 3: Cross-Embodiment Vision Embeddings (Left): We use t-SNE to visualize cross-embodiment latent embeddings generated by robot and demonstrator when executing different tasks on all three datasets. TCC Failure Example (Right): The robot and video clip are equivalent, but specific frames have high TCC losses. For example, a frame showing the robot performing the ’kettle’ action has a high loss due to its nearest neighbor in the video performing both ’kettle’ and ’light’ actions. This frame cycles back to the robot performing ’light’, which is mismatched.
Imagining Demonstration Videos. Algorithm 1 details the retrieval process for imaging a sequence of demonstration embeddings. We break the robot’s video into short-horizon windows and compare the embeddings with those from the play dataset, retrieving snippets with the lowest sequence-level distance. These retrieved snippets are concatenated to form an imagined long-horizon demonstration video. The key challenge is defining a distance function $d(\mathbf{z_{R}},\mathbf{z_{H}})$ that can handle video sequences of varying lengths. We propose two methods to compute this distance.
Method 1: Optimal Transport Distance. We calculate the Wasserstein distance (Optimal Transport) [47] between the robot and demonstrator video embeddings, i.e., the cost of the optimal transport plan that transfers one sequence of video embeddings into another. The robot’s embedding distribution is defined as $\rho_{R}=\{1/T,1/T,\dots,1/T\}$, and the demonstrator’s embedding distribution is defined as $\rho_{H}=\{1/T^{\prime},1/T^{\prime},\dots,1/T^{\prime}\}$, where $T$ and $T^{\prime}$ are the lengths of the video sequences respectively. The cost function for the transport is $C\in\mathbb{R}^{T\times T^{\prime}}$ where $C^{ij}$ is the cosine distance between the robot embedding $z_{R}^{i}$ and the demonstration embedding $z_{H}^{j}$. Our goal is to find the optimal assignment $M\in\mathbb{R}^{T\times T^{\prime}}$ that transports the distribution from $\rho_{R}$ to $\rho_{H}$ while minimizing the cost of the plan. Formally, we need to find $M^{*}=\operatorname*{arg\,min}\limits_{M}\sum_{i}\sum_{j}C^{ij}M^{ij}$. After solving the optimal transport assignment, the distance function is the cost of the plan, i.e., $d(\mathbf{z_{R}},\mathbf{z_{H}})=\sum_{i}\sum_{j}C^{ij}M^{*ij}$. In practice, we optimize an entropy-regularized version of this problem to find an approximate solution efficiently using the Sinkhorn-Knopp algorithm [47].
Method 2: Temporal-Cyclic Consistency (TCC) Distance. We calculate the TCC loss between human and robot videos following  [48] which computes cycle consistency between robot video embedding $\mathbf{z_{R}}=\{z_{R}^{1},z_{R}^{2},\dots,z_{R}^{T}\}$ and demonstrator $\mathbf{z_{H}}=\{z_{H}^{1},z_{H}^{2},\dots,z_{H}^{T^{\prime}}\}$. For each robot frame $z_{R}^{t}$, we first compute a similarity distribution $\mathbf{\alpha}$ of $z^{R}_{t}$ with respect to the demonstrator’s embeddings, to find a soft-nearest neighbor $\tilde{z}_{H}=\sum_{t^{\prime}=1}^{T^{\prime}}\alpha_{t}z^{t^{\prime}}_{H}$. Then, $\tilde{z}_{H}$ cycles back to the robot video by again computing its similarity distribution $\mathbf{\beta}$ with respect to robot video embeddings to get its soft-nearest neighbor $\tilde{z}_{R}^{t}=\sum_{t=1}^{T}\beta^{t}z^{t}_{R}$. The TCC distance for a robot frame $z_{R}^{t}$ is the mean square error with its cycled-back frame $\tilde{z}_{R}^{t}$ as $\ell_{\rm tcc}=||z_{R}^{t}-\tilde{z}_{R}^{t}||_{2}$. We define the video-level TCC distance function by summing over the frame-level losses $d(\mathbf{z_{R}},\mathbf{z_{H}})=\sum_{t=1}^{T}\ell_{\rm tcc}(z_{R}^{t})$.
We hypothesize that video retrieval using TCC distance can be inaccurate in two cases: (1) When robot and demonstrator embeddings differ due to variations in execution speed or style, leading to poor frame alignment. (2) When multiple robot embeddings correspond to a single demonstrator frame, as in sequential robot tasks versus parallel demonstrator actions, causing ambiguity in matching.
V Experiments
Simulation Setup and Datasets. We evaluate our approach using the Franka Kitchen simulator [49], where a 7-DOF Franka arm performs $7$ different tasks. We generate $3$ cross-embodiment video datasets each progressively increasing the embodiment and execution mismatch, which contain $580$ long-horizon ($\sim$25 seconds) robot trajectories completing a sequence of 4 tasks and a bank of demonstrator play data ($>$ 3 hours) for training our models. First, in Sphere-Easy, we replace the robot’s visual rendering with a sphere following the gripper’s position, creating a visual gap between robot and demonstrator. Second, in Sphere-Medium, we introduce manipulation style mismatches by applying randomized motion primitives to the demonstrator, such as the robot dragging an object while the demonstrator lifts and carries it. Finally, in Sphere-Hard, we create a further divergence where the demonstrator perform two tasks simultaneously, similar to how humans use two hands. Fig. 2 (left) illustrates the cross-embodiment datasets. Before training, all methods pre-train visual representations on unpaired robot and demonstrator videos, and policies are trained using Diffusion Policy[50].
Baselines. XSkill [4] simply conditions on robot videos during train-time, and uses its shared representation space to generalize zero-shot to inputs of demonstrator videos at test-time. OraclePairing [7, 8] is the gold-standard approach, assuming an oracle pairs demonstration with a robot trajectory, enabling conditioning on the demonstrator at train time. Our approach, RHyME, finds a middle ground. Without pairing, it imagines demonstrator trajectories that perform the same tasks as a robot trajectory (Section IV-B) by exploiting sequence-level correspondences. We compare two variants of our algorithm, RHyME-TCC and RHyME-OT, which differ in their distance functions used for retrieval. We also show how vision representations can be improved using short-horizon robot-demonstrator task pairs.
Evaluation and Success Metrics. At test-time, we provide the robot policy with demonstrator videos consisting of composition of 4 tasks not seen during training. We evaluate one-shot imitation performance across 20 different demonstrator videos, rolling out the robot policy from the last 5 model checkpoints, yielding 100 total trials per dataset. We measure Task Recall, which measures recall by counting the successfully completed tasks shown in the demonstrator’s video, and Task Imprecision, which measures imprecision and reports the percentage of tasks the robot attempts incorrectly—those not specified in the demonstrator’s video.
Q1. How does performance vary across different levels of execution mismatch?
As the demonstrator’s actions visually and physically deviate further from those of the robot, policies trained with our framework RHyME consistently outperform XSkill (Fig. 2). The OraclePairing baseline consistently achieves greater than $90\%$ task recall with minimal task imprecision. In the prescence of merely a cosmetic gap in Sphere-Easy, all methods have high task recall while maintaining precision. Still, we see that RHyME-OT has higher $90\%$ task recall compared to XSkill’s $82\%$. This gap increases on the Sphere-Medium dataset ($72\%$ vs $42\%$) and becomes larger on Sphere-Hard ($53\%$ vs $1\%$).
Fig. 3 (left) investigates this trend by probing the visual representations of the video encoder $\phi$, common across policies. We plot the image embeddings of the robot and demonstrator across three different tasks using a t-SNE plot. In Sphere-Easy, robot and demonstrator embeddings are closely clustered by task execution, so the XSkill policy can zero-shot transfer to demonstration embeddings at test time, but quickly drops in performance as the execution mismatch increases. RHyME algorithms overcome this problem and successfully retrieve the correct demonstration videos at train-time by reasoning over sequences of embeddings. The demonstrator performs two tasks simultaneously in Sphere-Hard. Consequently, demonstrator embeddings during two simultaneous tasks lie between the individual task embedding clusters of the robot. RHyME-OT maintains its performance relative to baseline as a result, but we observe that RHyME-TCC suffers a significant drop in task recall compared to RHyME-OT ($34\%$ vs $53\%$). More interestingly, this method has much higher task imprecision, even in the easier Sphere-Medium dataset. This suggests inaccurate retrieval of demonstration videos at train-time.
Figure 4: Optimal Transport Distances. We measure the similarity between robot and demonstrator videos on the Sphere-Hard dataset by computing the cost of the Optimal Transport (OT) plans. The sum over the entire transport cost matrix costs yields the distance between videos. OT costs are lowest when tasks are the same between videos (highlighted by a tick mark).
Q2. How does video retrieval using Optimal Transport and TCC impact policies at test-time?
When task embeddings are well clustered (Sphere-Easy, Fig. 3), TCC correctly cycles back and forth within clusters to yield low distance for videos matching in task specification, and high distance otherwise. However, as task embedding clusters deviate due to execution mismatches, we observe inaccuracies: in Sphere-Hard (Fig. 3 (right)) when both clips complete the same two tasks, a bimanual task embedding lies in between two robot task clusters which, resulting in cycling-back to the incorrect robot frame. We see that RHyME-TCC has relatively high task recall with lower task imprecision (due to incorrect retrievals) relative to other methods. RHyME-OT performs strictly better across datasets (Fig. 2).
The key reason for this performance difference is that optimal transport computes distances by matching videos across a sequence of embeddings. Fig. 4 visualizes the cost of the optimal transport plan between prompt robot clips and demonstration videos in the hard Sphere-Hard dataset. Comparing a robot clip doing two tasks (e.g. kettle and light), the transport cost across assignments is minimum only when compared to the demonstrator performing those same two tasks. TCC, on the other hand, attempts to establish one-to-one correspondences between the robot and demonstration frames, which are lacking in this dataset.
Figure 5: Performance Improves by Pairing Skills on Sphere-Hard. For both non-retrieval and retrieval-based methods, performance improves when fine-tuned when the visual encoder is finetuned short-horizon robot-demonstrator snippet pairs using a contrastive optimal transport loss.
Q3. Does fine-tuning visual representations with task-equivalent pairs improve one-shot imitation?
Following Section IV-A, we perform a study assuming access to short-horizon task pairings across embodiments to fine-tune our representation space with the contrastive optimal transport-based task alignment loss $L_{task}(\phi)$ in the Sphere-Hard setting. We find that encouraging induced distributions over embeddings to be similar lifts the performance of both XSkill and RHyME-OT (Fig. 5), and scales up as more data is accessible for fine-tuning since task alignment helps accurately retrieve video snippets for policy training. Ultimately, comparing induced distributions over embeddings with optimal transport is a beneficial design choice for matching clips at a task-level in the face of execution mismatches, as RHyME-OT (0% fine-tuned) still significantly outperforms XSkill (fine-tuned).
VI Discussion and Limitations
This work addresses the challenge of one-shot imitation in the presence of mismatched execution by the demonstrator. We propose RHyME, a novel framework that leverages task-level correspondences to bridge frame-level visual disparities between the robot and the demonstrator, enabling the learning of a video-conditioned policy without paired data.
Limitations. While the exact test-time task compositions are unseen during training, our method relies on transitions between task pairs in the robot dataset to learn transition actions. This limits the ability to learn entirely new task sequences. We note still that our method generalizes well to new compositions when such transitions are present. Expanding the dataset could enhance its generalizability.